# Infrequent Resolving Algorithm for Online Linear Programming

### v3

@@ Abstract
Online linear programming (OLP) has gained significant attention from both researchers and practitioners due to its extensive applications, such as online auction, network revenue management and advertising. Existing OLP algorithms fall into two categories: LP-based algorithms and LP-free algorithms. The former one typically guarantees better performance, even offering a constant regret, but requires solving a large number of LPs, which could be computationally expensive. In contrast, LP-free algorithm only requires first-order computations but induces a worse performance, lacking a constant regret bound. In this work, we study the case where the inputs are drawn from an unknown finite-support distribution, and bridge the gap between these two extremes by proposing an algorithm that achieves a constant regret while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$. Moreover, when we are allowed to solve LPs only $M$ times, we propose an algorithm that can guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret. Furthermore, when the arrival probabilities are known at the beginning, our algorithm can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times. Numerical experiments are conducted to demonstrate the efficiency of the proposed algorithms.
@@ 1 Introduction
Online linear programming (OLP) is a classical problem in online decision making. In this problem, a decision maker manages multiple types of resources, such as airplane seats or hotel rooms, with limited inventory. Customers arrive sequentially, each requesting a combination of resources (e.g., multi-leg flights or multi-day room reservations) and offering a bid price. Upon observing an arriving customer’s request and bid price, the decision maker needs to make an irrevocable decision to accept or reject the request in order to maximize the total expected revenue without violating resource constraints. OLP has applications in various industries, including online auction (Buchbinder et al. 2007), advertisement (Mehta et al. 2005), covering and packing (Buchbinder and Naor 2009b), e-commerce fulfillment (Jasin and Sinha 2015), and network revenue management (Jasin 2015), among many others. To facilitate our discussion, we first provide the formal formulation of the OLP problem.
Consider a decision maker with $m$ types of resources facing sequentially arriving customers over $T$ discrete time periods. We assume the initial inventory is $T\boldsymbol{\rho}$, where $\boldsymbol{\rho}=(\rho_{1},\dots,\rho_{m})\in\mathbb{R}_{+}^{m}$ is given and denotes the vector of average inventories per period. Each customer arriving at period $t$ is characterized by a reward $r_{t}\in\mathbb{R}_{+}$ and a consumption vector $\boldsymbol{A}_{t}\in\mathbb{R}_{+}^{m}$. The time periods are indexed forward, starting from period $t=1$ and ending at period $t=T$. In each time period $t$, after observing the arriving customer characterized by $(r_{t},\boldsymbol{A}_{t})$, the decision maker must irrevocably decide whether to accept or reject the customer ($x_{t}\in\{0,1\}$) [^1] Our results can be easily extended to the case with multiple actions. without knowing future information. Given full information $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$, we can solve the following (offline) linear program:
@@
In this paper, we assume the customers’ features $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$ with $n\in\mathbb{Z}_{+}$ supports. Such an assumption is widely adopted in the literature (e.g., Jasin 2015, Xie et al. 2023 and Chen et al. 2024). With this assumption, the OLP problem can be simplified as follows:
There are $n$ types of arrivals with corresponding coefficient groups $\{(r_{j},\boldsymbol{A}_{j})\}_{j=1}^{n}$, and each arrival’s type is drawn according to the distribution $\{p_{j}\}_{j=1}^{n}$ (with $\sum_{j=1}^{n}p_{j}=1$). Slightly abusing the notations, we define $\boldsymbol{r}\in\mathbb{R}_{+}^{n}$ and $\boldsymbol{A}\in\mathbb{R}_{+}^{m\times n}$ as the reward vector and the consumption matrix for all types of customers, respectively. Let $j^{t}$ denote the random type of the $t$-th arrival. The decision maker’s problem is to select a non-anticipative policy $\mu$, to maximize the total expected rewards while satisfying the inventory constraints. The OLP problem can be written as follows:
@@
where $\Pi$ is the set of all non-anticipative policies, $x^{t}_{\mu,j}$ denotes whether the policy $\mu$ accepts a type-$j$ customer at time $t$, and $Y_{j}^{t}\in\{0,1\}$ is the random indicator whether a type-$j$ customer arrives at time $t$. If we relax the integer constraints in (2) and replace all random variables with their expectations, then the model is referred to as the “fluid model”.
@@
Broadly speaking, OLP algorithms in prior studies can be categorized into two types: LP-based algorithms and LP-free algorithms. LP-based algorithms typically make decisions based on the optimal primal/dual solution to the fluid LP in each period, offering strong performance guarantees but requiring solving a large number of LPs (e.g., Li and Ye 2022, Xie et al. 2023, Chen et al. 2024, etc.). Although commercial solvers can efficiently solve LPs, frequent LP resolving remains computationally expensive especially for large-scale or time-sensitive problems. In response, researchers recently propose LP-free algorithms that typically use gradient descent methods to derive an approximate solution converging to the optimal dual solution to the fluid LP, and makes decisions based on the approximate solution. Thus, such algorithms rely only on first-order computations and avoid solving any LP (e.g., Li et al. 2020, Gao et al. 2024, Ma et al. 2024, etc.). However, these LP-free approaches usually lead to weaker performance bound than the LP-based algorithms.
@@
In this paper, we aim to strike a balance between the performance guarantee of LP-based algorithms and the computational efficiency of LP-free algorithms. Specifically, we propose an algorithm which achieves $\mathcal{O}(1)$ regret for the OLP problem while requiring solving LPs only $\mathcal{O}(\log\log T)$ times. In our algorithm, we concentrate the resolving (of an approximate LP) periods 1) on the beginning of the time horizon and 2) toward the end of the time horizon. The first set of resolvings aim to update the policy when the data is scarce so that correcting learning errors is important, while the latter set of resolvings aim to update the policy when the inventory is running out. Between two resolvings, we use the solution of an approximate LP to guide the allocation. Moreover, when we are allowed to solve LPs for up to $M$ times, we propose an algorithm that can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret. In addition, we show that the proposed algorithms can be easily adapted to the settings where the arrival probabilities of each type of request are known, which also lead to near-optimal regrets in those settings. In this case, a modified algorithm can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times. Therefore, our results demonstrate that one can achieve near optimal regrets only with a few resolvings, and depict the precise tradeoff between the frequency of resolving and the performance of the algorithm.
@@
We would like to highlight the technical contribution of this work. In the literature of LP-based algorithms for solving OLP, the algorithm usually solves an updated fluid model in each time period (e.g., Vera and Banerjee 2021, Xie et al. 2023, Chen et al. 2024). This approach allows bridging their policies with the optimal hindsight policy through two LPs: the fluid model and the hindsight benchmark. In contrast, our approach solves the updated fluid model only at a few selected periods ($\mathcal{O}(\log\log T)$ periods), and hence cannot access the optimal solution of the updated fluid model in most periods, posing a challenge for the analysis. In order to overcome this technical challenge, we approximate the optimal solution of the updated fluid model based on the latest obtained fluid solution and some first-order computations. While this solution is sub-optimal for the updated fluid model, we prove that it is optimal to a surrogate LP with high probability. Therefore, we can bridge our policy with the optimal hindsight policy through the surrogate LP and the hindsight LP, from which we can obtain the desired result.
@@ 1.1 Literature Review
Unknown Distribution. We start with the stream assuming unknown arrival probabilities. Our work closely relates to studies under the random input assumption, where the coefficients $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$. We focus on the asymptotic order of the regret as $T$ increases. Bray (2024) shows that the regret is at least $\Omega(\log T)$ when the demand distribution has infinite supports. In this paper, we focus on the case with a finite-support demand distribution. We first review the so called LP-based algorithms which require solving many LPs but guarantee good performance. For example, under certain non-degeneracy assumptions, Jasin (2015) proposes an $\mathcal{O}(\log^{2}T)$-regret algorithm that requires solving LPs $\mathcal{O}(\log T)$ times. Subsequently, Chen et al. (2024) consider a similar algorithm that requires solving LPs $T$ times, achieving $\mathcal{O}(1)$ regret under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T}\log T)$ otherwise. Under the non-degeneracy assumption and some distribution assumptions, Li and Ye (2022) provide an $\mathcal{O}(\log T\log\log T)$-regret algorithm from the dual perspective, and the algorithm requires solving LPs $T$ times. Then, under similar assumptions, Ma et al. (2024) provide an $\mathcal{O}(\log T)$-regret algorithm requiring solving LPs $\mathcal{O}(T)$ times. Recently, Xie et al. (2023) remove the non-degeneracy assumption and provide an OLP algorithm with $\mathcal{O}(1)$ regret but still requiring solving LPs $T$ times. Moreover, there are studies on infrequent resolving algorithms under additional assumptions. Given a complete sample path (i.e., $T$ samples), Freund and Banerjee (2019) propose an algorithm with lazy resolving (i.e., $\mathcal{O}(\log\log T)$ expected LP solving times) and prove a constant regret. Given a starting point with an accuracy of $\mathcal{O}(1/T)$, Ma et al. (2024) propose an infrequent resolving algorithm that can achieve $\mathcal{O}(\log T)$ regret under the non-degeneracy assumption by solving LPs $O(\log T)$ times. Both assumptions are stringent when $T$ is large. In this work, without such assumptions, we propose an algorithm that can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times.
Table 1: Comparison among OLP Algorithms (i.e., Unknown-Probability Case).
Paper
Regret
(Non-degenerate Case)
Regret
(Degenerate Case)
# of Resolvings
Jasin (2015) $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\log T)$
Li and Ye (2022) $\mathcal{O}(\log T\log\log T)$ $\mathcal{O}(T)$
Chen et al. (2024) $\mathcal{O}(1)$ $\mathcal{O}(\sqrt{T}\log T)$ $\mathcal{O}(T)$
Xie et al. (2023) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Ma et al. (2024) $\mathcal{O}(\log T)$ $\mathcal{O}(T)$
Freund and Banerjee (2019)
(Requiring $T$ samples )
$\mathcal{O}(1)$ $\mathcal{O}(1)$
$\mathcal{O}(\log\log T)$
(In expectation )
Ma et al. (2024)
(Requiring $\mathcal{O}(1/T)$-accuracy starting point )
$\mathcal{O}(\log T)$ $\mathcal{O}(\log T)$
Li et al. (2020) $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(\sqrt{T})$ 00
Gao et al. (2024) $\mathcal{O}(T^{1/3})$ 00
Ma et al. (2024) $\mathcal{O}(\log^{2}T)$ 00
This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $M$
@@
Although LP-based algorithms have good performance guarantees, frequent LP solving can be computational expensive for large-scale or time-sensitive problems. To address these computational challenges, recent studies have developed LP-free algorithms that never solve full LPs. Most of these studies analyze the regret in absence of budget constraints and derive bounds for both the regret and the constraint violation, i.e., the cumulative over-budget consumption. [^2] As Li et al. (2020) note, although an LP-free algorithm can be modified to a feasible policy satisfying the budget constraints, deriving a regret bound for such a modified policy remains an open question. The pioneering work Li et al. (2020) adopts a stochastic sub-gradient descent method to learn the optimal dual prices and provides an $\mathcal{O}(\sqrt{T})$-regret LP-free algorithm with $\mathcal{O}(\sqrt{T})$ constraint violation. This algorithm only requires first-order computations and never solves any full LP. Subsequently, researchers have studied variants of this algorithm and derived tighter bounds under the non-degeneracy assumption and the distribution assumptions in Li and Ye (2022). For example, Gao et al. (2024) propose variants that improve the bounds for both regret and the constraint violation to $\mathcal{O}(T^{1/3})$. Similarly, Ma et al. (2024) propose an LP-free algorithm which guarantees an $\mathcal{O}(\log^{2}T)$ regret without constraint violation. In this work, we achieve a balance between computational efficiency and algorithm performance by proposing an algorithm that achieves a constant regret bound by solving LPs $\mathcal{O}(\log\log T)$ times. A summary of the existing results and our result is presented in Table 1.
@@
Another stream of research studies the OLP problem under the random permutation assumption. Here, the set $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$ is adversarially chosen, but the arrival order is uniformly distributed over all the permutations. Researchers aim to derive necessary conditions for the existence of a $(1-\epsilon)$-competitive algorithm, see, e.g., Agrawal et al. (2014), Molinaro and Ravi (2014), Gupta and Molinaro (2014), Kesselheim et al. (2014). Since the concentration bound under the random permutation assumption is weaker, the lower bound in Agrawal et al. (2014) implies that the asymptotic order of the regret is $\Omega(\sqrt{T})$. We also highlight that Agrawal et al. (2014), Molinaro and Ravi (2014) and Gupta and Molinaro (2014) solve LPs $\mathcal{O}(\log T)$ times and their resolving periods concentrate at the beginning, sharing similarities with our resolving schedule.

### v4

@@ Abstract
Online linear programming (OLP) has gained significant attention from both researchers and practitioners due to its extensive applications, such as online auction, network revenue management, order fulfillment and advertising. Existing OLP algorithms fall into two categories: LP-based algorithms and LP-free algorithms. The former one typically guarantees better performance, even offering a constant regret, but requires solving a large number of LPs, which could be computationally expensive. In contrast, LP-free algorithm only requires first-order computations but induces a worse performance, lacking a constant regret bound. In this work, we bridge the gap between these two extremes by proposing a well-performing algorithm, that solves LPs at a few selected time points and conducts first-order computations at other time points. Specifically, for the case where the inputs are drawn from an unknown finite-support distribution, the proposed algorithm achieves a constant regret (even for the hard “degenerate” case) while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$. Moreover, when we are allowed to solve LPs only $M$ times, we design the corresponding schedule such that the proposed algorithm can guarantee a nearly $\mathcal{O}\left(T^{(1/2)^{M-1}}\right)$ regret. Our work highlights the value of resolving both at the beginning and the end of the selling horizon, and provides a novel framework to prove the performance guarantee of the proposed policy under different infrequent resolving schedules. Furthermore, when the arrival probabilities are known at the beginning, our algorithm can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and a nearly $\mathcal{O}\left(T^{(1/2)^{M}}\right)$ regret by solving LPs only $M$ times. Numerical experiments are conducted to demonstrate the efficiency of the proposed algorithms.
@@ 1 Introduction
Online linear programming (OLP) is a classical problem in online decision making. In this problem, a decision maker manages multiple types of resources, such as airplane seats or products, with limited inventory. Customers arrive sequentially, each requesting a combination of resources (e.g., multi-leg flights or bundled products) and offering a bid price. Upon observing an arriving customer’s request and bid price, the decision maker needs to make an irrevocable decision to accept or reject the request in order to maximize the total expected revenue without violating resource constraints. OLP has applications in various areas, including online auction (Buchbinder et al. 2007), advertisement (Mehta et al. 2005), covering and packing (Buchbinder and Naor 2009b), e-commerce fulfillment (Jasin and Sinha 2015), and network revenue management (Jasin 2015), among many others. To facilitate our discussion, we first provide the formal formulation of the OLP problem.
Consider a decision maker with $m$ types of resources facing sequentially arriving customers over $T$ discrete time periods. We assume the initial inventory is $T\boldsymbol{\rho}$, where $\boldsymbol{\rho}=(\rho_{1},\dots,\rho_{m})\in\mathbb{R}_{+}^{m}$ is given and denotes the vector of average inventories per period. Each customer arriving at period $t$ is characterized by a reward $r_{t}\in\mathbb{R}_{+}$ and a consumption vector $\boldsymbol{A}_{t}\in\mathbb{R}_{+}^{m}$. The time periods are indexed forward, starting from period $t=1$ and ending at period $t=T$. In each time period $t$, after observing the arriving customer characterized by $(r_{t},\boldsymbol{A}_{t})$, the decision maker must irrevocably decide whether to accept or reject the customer ($x_{t}\in\{0,1\}$) without knowing future information. Given full information $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$, we can solve the following (offline) integer linear program:
@@
In this paper, we assume that the customers’ features $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$ with $n\in\mathbb{Z}_{+}$ supports. Such an assumption is widely adopted in the online resource allocation literature (e.g., Jasin 2015, Xie et al. 2023 and Gupta 2024). It is also motivated by practical considerations. In many industries including airline and e-commerce, the resource combinations and the prices are designed by the seller, and only customers interested in one of these options will be considered. Thus, the number of customer types is the same as the number of provided options, which is typically finite. Moreover, we assume that the time horizon $T$ is known in advance. For example, in the airline industry, the time horizon is determined by the known departure dates of the flights (see, e.g., Talluri and Van Ryzin 1998, Jasin 2015, Chen et al. 2024). Without the knowledge of the time horizon, it will be hard to derive a policy with a good performance guarantee because the remaining time plays a crucial role in the decision process (see, e.g., Agrawal et al. 2014, Jasin and Kumar 2012 and Bumpensanti and Wang 2020). With the above assumptions, the OLP problem can be simplified as follows:
There are $n$ types of customers with type-$j$ customers characterized by the coefficients $(r_{j},\boldsymbol{A}_{j})$, and each arriving customer’s type is drawn according to a distribution $\{p_{j}\}_{j=1}^{n}$ (with $\sum_{j=1}^{n}p_{j}=1$). Slightly abusing the notations, we define $\boldsymbol{r}\in\mathbb{R}_{+}^{n}$ and $\boldsymbol{A}\in\mathbb{R}_{+}^{m\times n}$ as the reward vector and the consumption matrix for all types of customers, respectively. Let $j^{t}$ denote the random type of the $t$-th arrival. The decision maker’s problem is to select a non-anticipative policy $\mu$, to maximize the total expected rewards while satisfying the inventory constraints. The OLP problem can be written as follows:
@@
where $\Pi$ is the set of all non-anticipative policies, $x^{t}_{\mu,j}$ denotes whether the policy $\mu$ accepts a type-$j$ customer at time $t$, and $Y_{j}^{t}\in\{0,1\}$ is the random indicator whether a type-$j$ customer arrives at time $t$.
@@
Since (2) is computationally intractable, researchers typically design heuristic policies based on a relaxed problem. Specifically, if we relax the integer constraints in (2) and replace all random variables with their (estimated) expectations, then we can derive the following linear program (let $y_{j}$ denote the number of accepted type-$j$ customers, and $\hat{\boldsymbol{p}}^{t}$ denote the estimated probabilities at period $t$)
$\begin{split}\bar{V}(T):=\max_{\boldsymbol{y}\geq\boldsymbol{0}}\quad&\boldsymbol{r}^{\mathsf{T}}\boldsymbol{y}\\ \text{s.t.}\quad&\boldsymbol{A}\boldsymbol{y}\leq T\boldsymbol{\rho},\\ &\boldsymbol{y}\leq\hat{\boldsymbol{p}}^{1}\cdot T,\ \forall j,\end{split}$ (4)
which is referred to as the “fluid model” (at period 1). Given a period $t$, some algorithms may update the remaining inventory and the estimated future arrivals in (4), and resolve the (updated) fluid linear program.
@@
Broadly speaking, based on the number of LP resolvings, OLP algorithms in prior studies can be categorized into two types: LP-based algorithms and LP-free algorithms. LP-based algorithms typically make decisions based on the optimal primal/dual solution to the updated fluid LP (4) in each period, offering strong performance guarantees but requiring solving a large number of LPs (e.g., Li and Ye 2022, Xie et al. 2023, Chen et al. 2024, etc.). Although commercial solvers can efficiently solve LPs, frequent LP resolving remains computationally expensive especially for large-scale or time-sensitive problems. For example, the leading hotel booking platform Booking.com received about 560 million visits per month during 2022-2024, implying about 5 milliseconds average inter-arrival time (Statista 2024). In contrast, it typically consumes from seconds to minutes to solve linear programs of practical problems (see, e.g., Mittelmann 2024). Moreover, the optimal basis may change during the time horizon, and hence for each period we cannot simply use the last-period optimal solution to significantly speed up LP solving. In response, researchers recently propose LP-free algorithms that typically use gradient descent methods to derive an approximate solution converging to the optimal dual solution to the fluid LP, and makes decisions based on the approximate solution. Thus, such algorithms rely only on first-order computations and avoid solving any LP (e.g., Li et al. 2020, Gao et al. 2024, Ma et al. 2024, etc.). However, these LP-free approaches usually lead to weaker performance bound than the LP-based algorithms.
@@
In this paper, we aim to strike a balance between the performance guarantee of LP-based algorithms and the computational efficiency of LP-free algorithms. Specifically, we propose an algorithm which achieves $\mathcal{O}(1)$ regret for the OLP problem while requiring solving LPs only $\mathcal{O}(\log\log T)$ times. In our algorithm, we concentrate the resolving (of the fluid LP) periods 1) on the beginning of the time horizon and 2) toward the end of the time horizon. The first set of resolvings aim to update the policy when the data is scarce so that correcting learning errors is important, while the latter set of resolvings aim to update the policy when the inventory is running out so that the optimal policy varies dramastically. Between two resolvings, we use the optimal fluid solution of the latest resolving plus some first-order computations to guide the allocation. Moreover, when we are allowed to solve LPs for up to $M$ times, we propose an algorithm that can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret, where $\epsilon$ can be any positive constant. If we take $\epsilon\to 0$, the regret bound is about $\mathcal{O}\left(T^{(1/2)^{M-1}}\right)$. In addition, we show that the proposed algorithms can be easily adapted to the settings where the arrival probabilities of each type of request are known, which also lead to near-optimal regrets in those settings. In this case, a modified algorithm can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times. Therefore, our results demonstrate that one can achieve near optimal regrets only with a few resolvings, and depict the precise tradeoff between the frequency of resolving and the performance of the proposed algorithm.
In addition to the strong performance of our proposed algorithm, we would like to highlight one especially significant feature of our algorithm. In many previous studies, an important “non-degeneracy” assumption is imposed on the underlying input data and such an assumption would greatly affect the performance of proposed algorithms. Particularly, the “non-degeneracy” assumption refers to the assumption that the fluid model (4) with the estimation $\hat{\boldsymbol{p}}^{1}$ replaced by the true value $\boldsymbol{p}$ (called the “no-learning fluid model”) is non-degenerate. In some works (e.g., Gupta 2024, Wei et al. 2023), a $\delta$ is defined to be a measure of the distance between the current inventory configuration $T\boldsymbol{\rho}$ and the nearest inventory configuration under which the no-learning fluid model is degenerate, and the regret result is dependent on $1/\delta$ (thus their regret bounds dramastically increase and tend to infinity as $\delta$ tends to zero). Importantly, our results do not rely on the non-degeneracy assumption. In fact, to the best of our knowledge, our algorithm is the first that achieves a constant regret with such few resolvings for the case without the non-degeneracy assumption and distribution knowledge. To distinguish different results, we use $\mathcal{O}_{\delta}(\cdot)$ ($\mathcal{O}(\cdot)$, resp.) to denote regret bounds containing (without, resp.) $1/\delta$.
@@
Finally, we would like to highlight the technical contribution of this work. In the literature of LP-based algorithms for solving OLP, the algorithm usually solves an updated fluid model in each time period (e.g., Vera and Banerjee 2020, Xie et al. 2023, Chen et al. 2024). This approach allows updating the probability estimation frequently and bridging their policies with the optimal hindsight policy through two LPs: the fluid model and the hindsight benchmark. In contrast, our approach solves the updated fluid model only at a few selected periods ($\mathcal{O}(\log\log T)$ periods), and hence cannot update the probability estimation and access the optimal solution of the updated fluid model in most periods, posing challenges for the analysis. In order to overcome these technical challenges, we approximate the optimal solution of the updated fluid model based on the latest obtained fluid solution and some first-order computations. While this solution is sub-optimal for the updated fluid model for the corresponding period, we prove that it is optimal to a surrogate LP with high probability. Therefore, we can bridge our policy with the optimal hindsight policy through the surrogate LP and the hindsight LP, from which we can obtain the desired result.
@@ 1.1 Literature Review
Unknown Distribution. We start with the stream assuming unknown type distribution. Our work closely relates to studies under the random input assumption, where the coefficients $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$. For the underlying distribution, there are two diverging assumptions: finite-support distribution and continuous-support distribution.
@@
We start with the finite-support distribution assumption, which is the same as our setting. Under such an assumption, the distribution $\mathcal{P}$ is supported by finite bounded points such that arrivals can be categorized into finite types. We first review the so called LP-based algorithms which require solving many LPs but guarantee good performance. For example, under the non-degeneracy assumption (i.e., $\delta>0$), Jasin (2015) proposes an $\mathcal{O}_{\delta}(\log^{2}T)$-regret algorithm that requires solving LPs $\mathcal{O}(\log T)$ times. Subsequently, Chen et al. (2024) consider a similar algorithm that requires solving LPs $T$ times, achieving $\mathcal{O}_{\delta}(1)$ regret under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T}\log T)$ otherwise. Wei et al. (2023) propose a primal-dual algorithm that solves LPs $T$ times and the corresponding regret is $\mathcal{O}_{\delta}(1)$ under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T})$ in general. Xie et al. (2023) remove the non-degeneracy assumption and provide an OLP algorithm with $\mathcal{O}(1)$ regret but still requiring solving LPs $T$ times. As mentioned, although LP-based algorithms have good performance guarantees, frequent LP solving can be computational expensive for large-scale or time-sensitive problems. To address these computational challenges, recent studies have developed LP-free algorithms without LP resolving. The pioneering works Balseiro et al. (2020) and Li et al. (2020) adopt stochastic gradient descent methods to learn the optimal dual prices and provide $\mathcal{O}(\sqrt{T})$-regret LP-free algorithms. [^1] The algorithm in Li et al. (2020) also allows consuming $\mathcal{O}(\sqrt{T})$ additional budget. As Li et al. (2020) note, although the proposed LP-free algorithm can be modified to a feasible policy satisfying the budget constraints, deriving a regret bound for such a modified policy remains an open question. These algorithms only require first-order computations and never solve any full LP. In this work, we achieve a balance between computational efficiency and algorithm performance by proposing an algorithm that achieves a constant regret bound by solving LPs $\mathcal{O}(\log\log T)$ times. Moreover, we also provide regret bounds when the number of resolvings is finite, which to the best of our knowledge has not been provided in previous literature.
Table 1: Comparison among OLP Algorithms for Unknown-Probability Case.
Paper Regret (Non-degenerate Case) Regret (Degenerate Case) # of Resolvings Distribution Assumption Jasin (2015) $\mathcal{O}_{\delta}(\log^{2}T)$ $\mathcal{O}(\log T)$ Finite Chen et al. (2024) $\mathcal{O}_{\delta}(1)$ $\mathcal{O}(\sqrt{T}\log T)$ $\mathcal{O}(T)$ Finite Wei et al. (2023) $\mathcal{O}_{\delta}(1)$ $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(T)$ Finite Xie et al. (2023) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$ Finite Li and Ye (2022) $\mathcal{O}_{\delta}(\log T\log\log T)$ $\mathcal{O}(T)$ Continuous Bray (2024) $\mathcal{O}_{\delta}(\log T)$ $\mathcal{O}(T)$ Continuous Ma et al. (2024) $\mathcal{O}_{\delta}(\log T)$ $\mathcal{O}(T)$ Continuous Balseiro et al. (2020) $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(\sqrt{T})$ 00 General Li et al. (2020) $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(\sqrt{T})$ 00 General Gao et al. (2024) $\mathcal{O}_{\delta}(T^{1/3})$ 00 Continuous Ma et al. (2024) $\mathcal{O}_{\delta}(\log^{2}T)$ 00 Continuous This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ Finite This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $M$ Finite
Then, we review papers with the continuous-support distribution assumption. In this case, the distribution $\mathcal{P}$ is assumed to be supported by a bounded and continuous set and the probability density function is both lower bounded (away from 0) and upper bounded. (Note that finite-support distributions do not satisfy the assumption due to the existence of mass points. ) We still start with LP-based algorithms. For example, under the non-degeneracy assumption, Li and Ye (2022) provide an $\mathcal{O}_{\delta}(\log T\log\log T)$-regret algorithm from the dual perspective, and the algorithm requires solving LPs $T$ times. Then, Bray (2024) shows that the best possible regret bound for this problem is $\Omega(\log T)$. In addition, Bray (2024) and Ma et al. (2024) prove that the regret bound of Li and Ye (2022)’s algorithm is $\mathcal{O}_{\delta}(\log T)$ under the non-degeneracy assumption. Then, we introduce LP-free algorithms for this setting. The $\mathcal{O}(\sqrt{T})$ regret bounds established in Balseiro et al. (2020) and Li et al. (2020) also hold in this case. Subsequently, researchers have studied variants of this algorithm and tried to derive tighter bounds under the non-degeneracy assumption. For example, Gao et al. (2024) propose variants that improve the bounds to $\mathcal{O}_{\delta}(T^{1/3})$. [^2] The algorithm in Gao et al. (2024) also allows consuming $\mathcal{O}(T^{1/3})$ additional budget. Then, Ma et al. (2024) propose an LP-free algorithm which guarantees an $\mathcal{O}_{\delta}(\log^{2}T)$ regret. A summary of the existing results and our result is presented in Table 1 (where $\epsilon$ can be any positive constant). As mentioned, the regret bounds of some works are infinite under the degenerate case (i.e., $\delta=0$). We will use ‘-’ to denote such results.
Moreover, there are some works extending the previous settings. For example, for any bounded distribution (subsuming previous two classes), Balseiro et al. (2020), Li et al. (2020), Gao et al. (2023) and Balseiro et al. (2022) provide LP-free algorithms with $\mathcal{O}(\sqrt{T})$ regrets. Jiang et al. (2020) and Balseiro et al. (2023) consider the case when the underlying distribution is non-stationary. Besbes and Zeevi (2012) and Ferreira et al. (2018) consider the pricing problem without the knowledge of the demand function.
@@
We also note another stream of research studies under the random permutation assumption. Here, the set $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$ is adversarially chosen, but the arrival order is uniformly distributed over all the permutations. Researchers aim to derive necessary conditions for the existence of a $(1-\epsilon)$-competitive algorithm, see, e.g., Agrawal et al. (2014), Molinaro and Ravi (2014), Gupta and Molinaro (2014), Kesselheim et al. (2014). Since the concentration bound under the random permutation assumption is weaker, the lower bound in Agrawal et al. (2014) implies that the asymptotic order of the regret is $\Omega(\sqrt{T})$. We also highlight that Agrawal et al. (2014), Molinaro and Ravi (2014) and Gupta and Molinaro (2014) solve LPs $\mathcal{O}(\log T)$ times and their resolving periods concentrate at the beginning, sharing similarities with the first half of our resolving schedule.